For a one-tailed test (upper tail), a sample size of 30 at 99% confidence, t = 2.46 1.316 -2.740 0.344
Answer #1

Solution:

We have to obtain critical t value one tailed (upper tail) for a sample of size 30 at 99% confidence level.

If sample size is 30, then degrees of freedom = (30 - 1) = 29.

Confidence level = 99% = 0.99

Hence, significance level = (1- 0.99) = 0.01

Using t table, we can get critical t value (upper tailed) at 0.01 significance level and 29 degrees of freedom, which will be as follows:

t = t(0.01, 29) = 2.46 (upper tailed)

Hence, 1st option is correct.

Please rate the answer. Thank you.